OpenStudy (anonymous):

Use the change of base formula to evaluate log (base of five) 100. The convert it to a logarithm in base of three. Round to the nearest thousandth.

OpenStudy (anonymous):

\[\log_5(100)=\frac{\log(100)}{\log(5)}=\frac{2}{\log(5)}\]

OpenStudy (anonymous):

then a calculator